# Generalized Bayesian Multidimensional Scaling and Model Comparison

### v2

@@ 2.1.3 GBMDS with truncated Student’s $t$-distribution
To further relax the assumption of constant Gaussian error variance in the dissimilarity, we introduce the model with truncated Student’s $t$-distribution. Consequently, we can accommodate different degrees of uncertainty associated with dissimilarity by using different error variances. The $t$-distribution is often used as an alternative to the Gaussian distribution as a more robust model to fit data with heavier tails (Lange et al., 1989; Lin and Fong, 2019). In many applications, the outliers add more uncertainty around the tails of the dissimilarity measures. Fitting the truncated $t$-distribution provides a longer tail.
@@
[…]
@@
For a given matrix of dissimilarities D, the likelihood function of the latent variables $\textbf{X}=\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\right\}$, unknown parameters $\sigma^{2}$, $\zeta_{i,j}$ and $U$ under $\mathcal{M}_{TT}$, can be written as:
$l\left(\textbf{D}|\textbf{X},\sigma^{2},\zeta_{i,j},U,\mathcal{M}_{TT}\right)$ $\propto\left(\sigma^{2}\right)^{-\frac{m}{2}}\times$
$\exp\left\{\frac{1}{2}\sum_{\begin{subarray}{c}i,j=1\\ i>j\end{subarray}}^{n}\log\left(\zeta_{i,j}\right)-\frac{1}{2\sigma^{2}}\sum_{\begin{subarray}{c}i,j=1\\ i>j\end{subarray}}^{n}\zeta_{i,j}\left(d_{i,j}-\delta_{i,j}\right)^{2}\right.$
@@
where $\Phi(\cdot)$ and $m=n(n-1)/2$ are defined as in the model $\mathcal{M}_{TSN}$.
@@ 2.2.1 Prior distributions
Under the Bayesian framework, the prior distributions for the unknown parameters need to be specified in advance. We first introduce the prior distributions for model $\mathcal{M}_{TSN}$ with unknown parameters $\textbf{x}_{i}$, $\sigma^{2}$, and $\psi$. We assume prior independence among parameters. For the prior of $\textbf{x}_{i}$, we choose a multivariate Gaussian distribution with mean 0 and a diagonal covariance matrix $\Lambda=\text{diag}(\lambda_{1},\ldots,\lambda_{p})$. In other words, $\textbf{x}_{i}\sim\mathcal{N}\left(\textbf{0},\Lambda\right)$, independently for $i=1,\ldots,n$. For the elements along the diagonal covariance matrix, we assume an inverse Gamma distribution for the hyperprior distribution, i.e., $\lambda_{k}\sim\mathcal{IG}\left(\alpha,\beta_{k}\right)$, independently for $k=1,\ldots,p$. Under this assumption, we posit independence among the latent variables, resulting in a simpler model with reduced hyperparameters and improved interpretability. For $\sigma^{2}$, we use an inverse Gamma distribution for the prior distribution, i.e., $\sigma^{2}\sim\mathcal{IG}\left(a,b\right)$. For $\psi$, we choose a diffuse prior, i.e., $\psi\sim\mathcal{U}\left(c,d\right)$. We denote the prior distributions of the unknown parameters as $\pi\left(\textbf{X}|\Lambda\right)$, $\pi\left(\sigma^{2}\right)$, $\pi\left(\psi\right)$ and $\pi\left(\Lambda\right)$.
@@ 2.2.2 Posterior distributions
In the Bayesian framework, our interest is the posterior distribution on $\boldsymbol{s}$ given dissimilarity matrix D, denoted as:
$\pi\left(\boldsymbol{s}|\textbf{D}\right)=\frac{\gamma\left(\boldsymbol{s}|\textbf{D}\right)}{M}\propto l(\textbf{D}|\boldsymbol{s})\times\pi(\boldsymbol{s}),$ (7)
[…]
@@ 2.2.3 Adaptive Bayesian inference
Let $\textbf{X}^{(0)}$ be the hidden variables associated with objects $\textbf{Z}^{(0)}=\{\mathbf{z}_{1},\ldots,\mathbf{z}_{n_{0}}\}$, and $\textbf{X}^{(1)}$ be the hidden variables associated with objects $\textbf{Z}^{(1)}=\{\mathbf{z}_{n_{0}+1},\ldots,\mathbf{z}_{n_{0}+n_{1}}\}$. Here, $n_{0}$ refers to the dimension of the original observations, and $n_{1}$ refers to the incremental dimension of the additional observations. Given dissimilarity metric $\mathcal{D}$, dissimilarities $\textbf{D}^{(0)}$ is obtained from $\textbf{Z}^{(0)}$ and D is obtained from $\textbf{Z}=(\textbf{Z}^{(0)},\textbf{Z}^{(1)})$.
@@
$\pi\left(\textbf{X}^{(0)},\textbf{X}^{(1)},\boldsymbol{\theta}|\textbf{D}\right)\propto l\left(\textbf{D}|\textbf{X}^{(0)},\textbf{X}^{(1)},\boldsymbol{\theta}\right)\pi\left(\textbf{X}^{(0)}|\Lambda^{(0)}\right)\pi\left(\textbf{X}^{(1)}|\Lambda^{(1)}\right)\pi\left(\boldsymbol{\theta}\right).$ (8)
@@
When the previous dissimilarity matrix $\textbf{D}^{(0)}$ is not available, we denote $\textbf{D}^{(0)}=\emptyset$ and $\textbf{X}^{(0)}=\emptyset$. With our notation, the posterior distribution in Equation 7 is a special case of Equation 8 when $\textbf{D}^{(0)}=\emptyset$, $\textbf{X}^{(0)}=\emptyset$. Therefore, we will only focus on the adaptive Bayesian inference with Monte Carlo methods for Equation 8 in Section 3.
@@ 2.3 Model comparison
As described in the previous section, the function $g$ can take different forms. In most cases, the optimal form of $g$, the number of significant attributes $p$ or the applied dissimilarity metrics are unknown. In this section, we approach the problem of comparing a discrete set of Bayesian models with the Bayes factor. Consider two models $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ with different likelihoods and corresponding sets of parameters $\boldsymbol{s}_{1}$ and $\boldsymbol{s}_{2}$. In the context of this paper, $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ would correspond to two competing models. Examples of competing models could be $\mathcal{M}_{TSN}$ versus $\mathcal{M}_{TT}$, or one model under a different choice of dimension $p$. The Bayes factor is defined as the ratio of the posterior odds to the prior odds. When the two models have equal prior probability, the Bayes Factor reduces to the ratio of two marginal likelihood estimates, and is given by:
[…]
Bayes factor can provide support to either model; a Bayes factor greater than 1 indicates support for model 1 over model 2 and vice versa. A rule of thumb, as suggested in Kass and Raftery (1995), can be viewed as guidelines for model selection from a Bayes factor. According to these guidelines, a Bayes factor falling within the range of 1 to 3 shows weak evidence, while a Bayes factor spanning from 3 to 10 implies substantial evidence. A Bayes factor exceeding 10 is considered as strong evidence. A typical challenge for using the Bayes factor is the computation of the marginal likelihood estimates, especially for MCMC-based methods (Wang et al., 2020). Marginal likelihood estimation is not straightforward in MCMC-based methods, and additional sampling procedures are needed to obtain these estimates. Several methods have been proposed to address this issue (Chib and Jeliazkov, 2001; Skilling, 2004; Robert and Wraith, 2009), but each has its drawbacks. One additional limitation shared by all MCMC-based marginal likelihood estimators is that they are generally biased. In contrast, SMC naturally produces unbiased or nearly unbiased marginal likelihood estimators as a byproduct of sampling, enabling straightforward model comparison via the Bayes factor.
On the other side, in the frequentist view, STRESS is a commonly used measure of fit for the object configuration problem (Kruskal, 1964). STRESS value is defined as
@@
where $\hat{\delta}_{i,j}$ is the distance found from the estimated object configuration. MDS methods form an object configuration that minimizes the STRESS values. A smaller STRESS value indicates a better fit.
In this work, we will select the optimal Bayesian model and dimension $p$ using the marginal likelihood estimates. We will compare and evaluate the performances of CMDS and GBMDS using the STRESS value.
@@ 2.4 Identifiability in multidimensional scaling
Similar to other dimensional reduction methods, identification issues arise in the posterior inference of GBMDS. For instance, the center and direction of the estimated points can be arbitrary. To address this, we standardize all posterior samples of $\textbf{x}_{i}$’s using Procrustes transformations (Goodall, 1991). This method aligns configurations using a least-squares criterion by a combination of scaling, rotation, reflection, and translation. After the Procrustes transformation, we construct credible regions from the adjusted posterior samples of $\textbf{x}_{i}$ to quantify uncertainty.
@@ 3.1 Intermediate distributions and particle initialization
To conduct the Bayesian inference for the posterior distribution in Equation 8, we propose to design an artificial sequence of annealing intermediate target distributions following the ideas from the SMC literature (Neal, 2001; Del Moral et al., 2006, 2007; Wang et al., 2020). Specifically, we create a sequence of annealing intermediate target distributions $\{\pi_{r}(\boldsymbol{s})\}_{0\leq r\leq R}$, such that
$\pi_{r}\left(\boldsymbol{s}\right)\propto\gamma_{r}\left(\boldsymbol{s}\right)=\left(l\left(\textbf{D}|\boldsymbol{s}\right)\pi\left(\boldsymbol{s}\right)\right)^{\tau_{r}}\times\tilde{\pi}_{0}\left(\boldsymbol{s}\right)^{1-\tau_{r}},$ (9)
where $\tilde{\pi}_{0}(\boldsymbol{s})$ is a reference distribution that is generally easy to sample from (Fan et al., 2011), and $10=\tau_{0}<\tau_{1}<\cdots<\tau_{R}=$ 1 is a sequence of annealing parameters. If $\tau_{r}$ is zero, the distribution becomes the reference distribution $\tilde{\pi}_{0}(\boldsymbol{s})$. At the other extreme, the distribution is the posterior distribution of interest when the power $\tau_{r}$ equals 1.
@@
$\tilde{\pi}_{0}\left(\boldsymbol{s}\right)=\tilde{\pi}_{0}\left(\textbf{X}^{(0)}\right)\tilde{\pi}_{1}\left(\textbf{X}^{(1)}\right)\tilde{\pi}_{0}\left(\boldsymbol{\theta}\right).$ (10)
Preferably, the reference distributions should possess properties that allow for convenient sampling and proximity to the modes of the target distribution. For simplicity, we choose the reference distribution for $\boldsymbol{\theta}$ to be its prior distribution, i.e. $\tilde{\pi}_{0}(\boldsymbol{\theta})=\pi(\boldsymbol{\theta})$, and the reference distributions for $\textbf{X}^{(1)}$ to be its prior, $\tilde{\pi}_{1}\left(\textbf{X}^{(1)}\right)=\pi\left(\textbf{X}^{(1)}|\Lambda^{(1)}\right)$; the reference distributions for $\textbf{X}^{(0)}$, denoted $\tilde{\pi}_{0}\left(\textbf{X}^{(0)}\right)$, is set to be a Gaussian distribution.
@@
Input : (a) Dissimilarity: $^{(0)},\textbf{D}$ D; (b) Priors and reference distributions over $\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\right\}$ and model parameters $\boldsymbol{\theta}=\{\sigma^{2},\Lambda,\psi,\zeta\}$; (c) Number of particles: $K$.
2
3for $k\in\{1,2,\ldots,K\}$ do
@@
5             Initialize particles of $\textbf{X}^{(0)}_{0,k}\sim\tilde{\pi}_{0}(\textbf{X}^{(0)})$.
6      Initialize particles of $\textbf{X}^{(1)}_{0,k}\sim\tilde{\pi}_{1}(\textbf{X}^{(1)})$. Initialize particles of parameters with independent samples from prior distributions: $\{{\sigma^{2}_{0,k}},\Lambda_{0,k},\psi_{0,k},\zeta_{0,k}\}\sim\pi(\boldsymbol{\theta})$.
7

### v3

@@ 2.1.3 GBMDS with truncated Student’s $t$-distribution
To further relax the assumption of constant Gaussian error variance in the dissimilarity, we introduce the model with truncated Student’s $t$-distribution. Consequently, we can accommodate different degrees of uncertainty associated with dissimilarity by using different error variances. The $t$-distribution is often used as an alternative to the Gaussian distribution as a more robust model to fit data with heavier tails [lange1989robust, lin2019bayesian]. In many applications, the outliers add more uncertainty around the tails of the dissimilarity measures. Fitting the truncated $t$-distribution provides a longer tail.
@@
$t_{\nu}\left(x;\mu,\sigma^{2}\right)=\int_{0}^{\infty}\mathcal{N}\left(x;\mu,\frac{\sigma^{2}}{\zeta}\right)Gamma\left(\zeta;\frac{\nu}{2},\frac{\nu}{2}\right)d\zeta.$ (5)
@@
Under $\mathcal{M}_{TT}$, for a given matrix of dissimilarities D, the likelihood function $L$ with respect to X, $\sigma^{2}$, and $\zeta_{i,j}$, can be written as:
$L$ $\left(\textbf{X},\sigma^{2},\zeta_{i,j};\textbf{D},\mathcal{M}_{TT}\right)\propto\left(\sigma^{2}\right)^{-\frac{m}{2}}\exp\left\{\frac{1}{2}\sum_{\begin{subarray}{c}i,j=1\\ i>j\end{subarray}}^{n}\log\left(\zeta_{i,j}\right)-\frac{1}{2\sigma^{2}}\sum_{\begin{subarray}{c}i,j=1\\ i>j\end{subarray}}^{n}\zeta_{i,j}\left(d_{i,j}-\delta_{i,j}\right)^{2}\right$.
@@
where $\Phi(\cdot)$ and m=n(n1)/2m=n(n-1)/2 are defined as in the model $\mathcal{M}_{TSN}$.
@@ 2.2.1 Prior distributions
Under the Bayesian framework, the prior distributions for the unknown parameters need to be specified in advance. We first introduce the prior distributions for model $\mathcal{M}_{TSN}$ with unknown parameters $\textbf{x}_{i}$, $\sigma^{2}$, and $\psi$. We assume prior independence among parameters. For the prior of $\textbf{x}_{i}$, we choose a multivariate Gaussian distribution with mean 0 and a diagonal covariance matrix $\Lambda=\text{diag}(\lambda_{1},\ldots,\lambda_{p})$. In other words, $\textbf{x}_{i}\sim\mathcal{N}\left(\textbf{0},\Lambda\right)$, independently for $i=1,\ldots,n$. For the elements along the diagonal covariance matrix, we assume an inverse Gamma distribution for the hyperprior distribution, i.e., $\lambda_{k}\sim\mathcal{IG}\left(\alpha,\beta_{k}\right)$, independently for $k=1,\ldots,p$. Under this assumption, we posit independence among the latent variables, resulting in a simpler model with reduced hyperparameters and improved interpretability. For $\sigma^{2}$, we use an inverse Gamma distribution for the prior distribution, i.e., $\sigma^{2}\sim\mathcal{IG}\left(a,b\right)$. For $\psi$, we choose a diffuse prior, i.e., $\psi\sim\mathcal{U}\left(c,d\right)$. We denote the prior distributions of the unknown parameters as $\pi\left(\textbf{X}|\Lambda\right)$, $\pi\left(\sigma^{2}\right)$, $\pi\left(\psi\right)$, and $\pi\left(\Lambda\right)$.
@@ 2.2.2 Posterior distributions
In the Bayesian framework, our interest lies in the posterior distribution of $\boldsymbol{s}$ given the dissimilarity matrix D, denoted as:
$\pi\left(\boldsymbol{s}|\textbf{D}\right)=\frac{\gamma\left(\boldsymbol{s}|\textbf{D}\right)}{Z}\propto L(\boldsymbol{s};\textbf{D})\,\pi(\boldsymbol{s}),$ (7)
where $\gamma(\boldsymbol{s}|\textbf{D})=L(\boldsymbol{s};\textbf{D})\pi(\boldsymbol{s})$ denotes the unnormalized posterior distribution, $L(\boldsymbol{s};\textbf{D})$ is the likelihood function, $\pi(\boldsymbol{s})$ is the prior on the parameters, and $Z=p(\textbf{D})=\int{\gamma(\boldsymbol{s}|\textbf{D})\,d\boldsymbol{s}}$ is the marginal likelihood. The likelihood functions are specified in Equations 2.1.2 and 2.1.3 for $\mathcal{M}_{TSN}$ and $\mathcal{M}_{TT}$, respectively. The prior distributions are described in the previous subsection. Since the normalizing constant $Z$ is intractable, we will use Monte Carlo methods to approximate the posterior distributions, which will be detailed in Section 3.
@@ 2.2.3 Adaptive Bayesian inference
Let $\textbf{X}^{(0)}$ be the hidden variables associated with $\textbf{Z}^{(0)}=\{\mathbf{z}_{1},\ldots,\mathbf{z}_{n_{0}}\}$, and $\textbf{X}^{(1)}$ be the hidden variables associated with $\textbf{Z}^{(1)}=\{\mathbf{z}_{n_{0}+1},\ldots,\mathbf{z}_{n_{0}+n_{1}}\}$. Here, $n_{0}$ refers to the dimension of the original observations, and $n_{1}$ refers to the incremental dimension of the additional observations. Given dissimilarity metric $\mathcal{D}$, dissimilarities $\textbf{D}^{(0)}$ is obtained from $\textbf{Z}^{(0)}$ and D is obtained from $\textbf{Z}=\textbf{Z}^{(0)}\cup\textbf{Z}^{(1)}$.
@@
$\pi\left(\textbf{X}^{(0)},\textbf{X}^{(1)},\boldsymbol{\theta}|\textbf{D}\right)\propto L\left(\textbf{X}^{(0)},\textbf{X}^{(1)},\boldsymbol{\theta};\textbf{D}\right)\pi\left(\textbf{X}^{(0)}|\Lambda^{(0)}\right)\pi\left(\textbf{X}^{(1)}|\Lambda^{(1)}\right)\pi\left(\boldsymbol{\theta}\right).$ (8)
@@
When the previous dissimilarity matrix $\textbf{D}^{(0)}$ is not available, we denote $\textbf{D}^{(0)}=\emptyset$ and $\textbf{X}^{(0)}=\emptyset$. With our notation, the posterior distribution in Equation 7 is a special case of Equation 8 when $\textbf{D}^{(0)}=\emptyset$ and $\textbf{X}^{(0)}=\emptyset$. Therefore, we will only focus on the adaptive Bayesian inference with Monte Carlo methods for Equation 8 in Section 3.
@@ 2.3 Model comparison
As described in the previous section, the function $g$ can take different forms. In most cases, the optimal form of $g$, the number of significant attributes $p$, and the suitable dissimilarity metrics are unknown. In this section, we approach the problem of comparing a discrete set of Bayesian models with the Bayes factor. Consider two models $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ with different likelihoods and corresponding sets of parameters $\boldsymbol{s}_{1}$ and $\boldsymbol{s}_{2}$. In the context of this paper, $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ would correspond to two competing models. Examples of competing models could be $\mathcal{M}_{TSN}$ versus $\mathcal{M}_{TT}$, or one model under a different choice of dimension $p$. The Bayes factor is defined as the ratio of the posterior odds to the prior odds. When the two models have equal prior probability, the Bayes factor reduces to the ratio of two marginal likelihoods, given by:
$\text{BF}_{12}=\frac{\int\pi(\boldsymbol{s}_{1}|\mathcal{M}_{1})L(\boldsymbol{s}_{1};\textbf{D},\mathcal{M}_{1})\,d\boldsymbol{s}_{1}}{\int\pi(\boldsymbol{s}_{2}|\mathcal{M}_{2})L(\boldsymbol{s}_{2};\textbf{D},\mathcal{M}_{2})\,d\boldsymbol{s}_{2}}=\frac{p\left(\textbf{D}|\mathcal{M}_{1}\right)}{p\left(\textbf{D}|\mathcal{M}_{2}\right)}.$
The Bayes factor quantifies the support for one model over another; a value greater than 1 supports Model 1, while a value less than 1 supports Model 2. Following the guidelines in kass1995bayes, a Bayes factor between 1 and 3 provides negligible evidence, while a value between 3 and 20 suggests positive evidence. A Bayes factor between 20 and 150 indicates strong evidence, and a value exceeding 150 is considered very strong evidence.
A typical challenge for using the Bayes factor is the computation of the marginal likelihood estimates, especially for MCMC-based methods. Marginal likelihood estimation is not straightforward in standard MCMC, often necessitating additional, complex sampling procedures or post-processing [chib2001marginal, skilling2004nested, robert2009computational]. Furthermore, many MCMC-based estimators, such as the harmonic mean, can suffer from infinite variance or significant finite-sample bias. In contrast, standard SMC algorithms produce unbiased marginal likelihood estimates naturally as a byproduct of the sequential sampling process [del2004feynman]. While adaptive SMC strategies may introduce negligible bias, they generally avoid the complex implementation hurdles associated with MCMC-based evidence estimation [zhou2016toward, wang2020annealed].
On the other side, in the frequentist view, STRESS is a commonly used measure of fit for the object configuration problem [kruskal1964multidimensional]. STRESS value is defined as
@@
where $\hat{\delta}_{i,j}$ is the distance found from the estimated object configuration. A smaller STRESS value indicates a better fit.
In this work, we select the optimal Bayesian model and dimension $p$ using marginal likelihood estimates. Furthermore, we compare the performance of CMDS and GBMDS using the STRESS value.
@@ 2.4 Identifiability in multidimensional scaling
Similar to other dimensional reduction methods, identification issues arise in the posterior inference of GBMDS. For instance, the center and direction of the estimated points can be arbitrary. To address this, we standardize all posterior samples of $\textbf{x}_{i}$’s using Procrustes transformations [goodall1991procrustes]. This method aligns configurations using a least-squares criterion by a combination of scaling, rotation, reflection, and translation. After the Procrustes transformation, we construct credible regions from the adjusted posterior samples of $\textbf{x}_{i}$ to quantify uncertainty.
@@ 3.1 Intermediate distributions and particle initialization
To conduct the Bayesian inference for the posterior distribution in Equation 8, we propose to design an artificial sequence of annealing intermediate target distributions following the ideas from the SMC literature [neal2001annealed, del2006sequential, Moral2007, wang2020annealed]. Specifically, we create a sequence of annealing intermediate target distributions $\{\pi_{r}(\boldsymbol{s})\}_{0\leq r\leq R}$, such that
$\pi_{r}\left(\boldsymbol{s}\right)\propto\gamma_{r}\left(\boldsymbol{s}\right)=\left(L\left(\boldsymbol{s};\textbf{D}\right)\pi\left(\boldsymbol{s}\right)\right)^{\tau_{r}}\times\tilde{\pi}\left(\boldsymbol{s}\right)^{1-\tau_{r}},$ (9)
where $\tilde{\pi}(\boldsymbol{s})$ is a reference distribution that is generally easy to sample from [fan2011choosing], and $0=\tau_{0}<\tau_{1}<\cdots<\tau_{R}=1$ is a sequence of annealing parameters. If $\tau_{r}$ is zero, the distribution becomes the reference distribution $\tilde{\pi}(\boldsymbol{s})$. At the other extreme, the distribution is the posterior distribution of interest when the power $\tau_{r}$ equals 1.
@@
$\tilde{\pi}\left(\boldsymbol{s}\right)=\tilde{\pi}_{0}\left(\textbf{X}^{(0)}\right)\tilde{\pi}_{1}\left(\textbf{X}^{(1)}\right)\tilde{\pi}_{\theta}\left(\boldsymbol{\theta}\right).$ (10)
Preferably, the reference distributions should possess properties that allow for convenient sampling and proximity to the modes of the target distribution. For simplicity, we choose the reference distribution for $\boldsymbol{\theta}$ to be its prior distribution, i.e. $\tilde{\pi}_{\theta}(\boldsymbol{\theta})=\pi(\boldsymbol{\theta})$, and the reference distributions for $\textbf{X}^{(1)}$ to be its prior, $\tilde{\pi}_{1}\left(\textbf{X}^{(1)}\right)=\pi\left(\textbf{X}^{(1)}|\Lambda^{(1)}\right)$; the reference distributions for $\textbf{X}^{(0)}$, denoted $\tilde{\pi}_{0}\left(\textbf{X}^{(0)}\right)$, is set to be a Gaussian distribution, detailed in Section 3.3.
@@
Input : (a) Dissimilarity $\textbf{D}^{(0)},\textbf{D}$; (b) Reference distributions $\tilde{\pi}_{0},\tilde{\pi}_{1},$ and $\tilde{\pi}_{\theta}$ for $\textbf{X}^{(0)},\textbf{X}^{(1)}$, and $\boldsymbol{\theta}=\{\sigma^{2},\Lambda,\psi,\zeta\}$, respectively. (c) Number of particles $K$.
@@
21exfor $k\in\{1,2,\ldots,K\}$ do
@@
4     Initialize particles of $\textbf{X}^{(0)}$: $\textbf{X}^{(0)}_{0,k}\sim\tilde{\pi}_{0}(\cdot)$.
5 Initialize particles of $\textbf{X}^{(1)}$: $\textbf{X}^{(1)}_{0,k}\sim\tilde{\pi}_{1}(\cdot)$.
6 1exInitialize particles of parameters: $\{{\sigma^{2}_{0,k}},\Lambda_{0,k},\psi_{0,k},\zeta_{0,k}\}\sim\tilde{\pi}_{\theta}(\cdot)$.
7 1exSet $\boldsymbol{s}_{0,k}\leftarrow\left\{\textbf{X}^{(0)}_{0,k},\textbf{X}^{(1)}_{0,k},\sigma^{2}_{0,k},\Lambda_{0,k},\psi_{0,k},\zeta_{0,k}\right\}$.